Imagineering, Inc., is considering an investment in CAD-CAM compatible design software with the cash flow profile shown in thWhat is the annual worth of this investment?
Answer #1

PW = -12000 - 1000 * (P/F, 18%,1) + 5000 * (P/A, 18%,6) *(P/F, 18%,1) - 3000 * (P/F, 18%,3) - 3000 * (P/F, 18%,6)

= -12000 - 1000 * 0.847458 + 5000 * 3.497603 * 0.847458 - 3000 * 0.608631 - 3000 * 0.370432

= -964.29

AW = -964.29 * (A/P, 18%,7)

= -964.29 * 0.262362

= -252.99